OpenStudy (anonymous):

Use the Distributive Property to find the price of 9 CDs that cost $13.99 each....I CHOSE THIS ANSWER: $125.91. Is it correct?

OpenStudy (texaschic101):

9(13 + 0.99) = 9 * 13 + 9 * 0.99 = 117 + 8.91 = 125.91 yep...thats right

OpenStudy (anonymous):

Oh, thank you very much :) Can I run a few more questions by you to confirm my answers? I also got 34/39 for this question. Is that correct? Teesha is in the French club. There are 39 students in the club. The French teacher will pick 5 students at random to guide visiting students from France. What is the probability that Teesha will not be picked as a guide?

OpenStudy (anonymous):

@texaschic101